# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

### v3

@@ Abstract
This paper investigates the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium (NE) in potential games. In such games, an efficient NE is defined as the maximizer of the potential function. Previous literature provides asymptotic convergence rates to efficient Nash equilibria, and existing finite-time rates are limited to potential games with further assumptions such as the interchangeability of players. In this paper, we prove the first finite-time convergence to an $\epsilon$-efficient NE in general potential games. Our bounds depend polynomially on $1/\epsilon$, an improvement over previous bounds that are exponential in $1/\epsilon$ and only hold for subclasses of potential games. We then strengthen our convergence result in two directions: first, we show that a variant of log-linear learning that requires a factor $A$ less feedback on the utility per round enjoys a similar convergence time; second, we demonstrate the robustness of our convergence guarantee if log-linear learning is subject to small perturbations such as alterations in the learning rule or noise-corrupted utilities.
keywords:
Game theory, potential games, log-linear learning, efficient Nash equilibrium.
@@ 1 Introduction
Interactions of multiple agents are at the heart of many applications in transportation networks, auctions, telecommunication networks, and multi-robot systems. In game theory, the Nash equilibrium is a popular solution concept to describe outcomes of a multi-agent system (Nash, 1951). Fundamental considerations in game theory are whether an NE exists, if strategic players can learn it, and if so at which speed they can learn it. Furthermore, for practical purposes it is important to understand which NE is learned. This is particularly pertinent in games that admit a social welfare function, as it enables the definition of an efficient NE as an NE that maximizes social welfare. The social welfare is typically an aggregate measure of individual utilities such as their sum or a measure of fairness. In distributed control of engineering systems, for example, studying efficient NEs is important, as the aim is to optimize some global objective function in a distributed manner.
The class of potential games (Monderer and Shapley, 1996b) lends itself to studying efficient NEs since every joint action maximizing the potential corresponds to an NE. This property follows from the fact that in potential games, the difference in a player’s utility generated by a unilateral change of her action equals the difference in potential. Consequently, if the social welfare is aligned with the potential function, meaning that an increase in social welfare is associated with an increase in potential (Paccagnan et al., 2022), then any joint action maximizing the potential is an efficient NE (Marden and Shamma, 2015). In identical interest games, for example, maximizing the aggregated utility is achieved by maximizing the potential function which is trivially given by the common utility function. In the example of coverage problems, the goal is to assign players to locations to achieve maximal coverage. When each player’s utility is designed as the marginal contribution then this goal is achieved by maximizing the resulting potential function (Marden and Wierman, 2013).
In this paper, we study the speed of convergence to approximate efficient NE in potential games.
@@ 1.1 Related work
In potential games, many learning rules are proven to converge to an NE, such as iterative best-response dynamics (Rosenthal, 1973; Awerbuch et al., 2008; Chien and Sinclair, 2011), no-regret algorithms (Krichene et al., 2015; Palaiopanos et al., 2017; Heliou et al., 2017), and fictitious play (Monderer and Shapley, 1996a, b). However, only log-linear learning (Blume, 1993; Young, 1993) is known to converge to an efficient Nash equilibrium. [^1] In log-linear learning, players asynchronously choose actions with probability proportional to their exponentiated utility. Due to this desirable property, variants of log-linear learning were proposed in subsequent works (Marden et al., 2007; Marden and Shamma, 2012). For example, Leslie and Marden (2011) proves that log-linear learning also handles the practical setting where the observed utilities are corrupted by noise. Moreover, Arslan et al. (2007) proposes binary log-linear learning, a slight modification of log-linear learning that only requires two points of feedback per round. Furthermore, (Tatarenko, 2017) provides an asymptotic rate of convergence to an efficient NE when the parameter controlling the players’ rationality in log-linear learning is time-varying. Another line of research studies how fast log-linear learning converges to its stationary distribution (Auletta et al., 2010, 2011, 2012) and how it behaves before reaching the stationary distribution (Auletta et al., 2012), but do not focus on Nash equilibria.
Few past works provide finite-time guarantees for log-linear learning to an $\epsilon$-efficient NE, an action profile whose potential is $\epsilon$-close to the maximum value. For instance, for atomic routing games with polynomial costs of degree at most $p$, Asadpour and Saberi (2009) proves a convergence time exponential in $p$ and $1/\epsilon$ and polynomial in $N$, where $N$ denotes the number of players. Moreover, Montanari and Saberi (2008, 2010) study games with graph structures between players and prove an exponential convergence time in $N$ and $1/\epsilon$ in the worst case. Finally, in potential games with interchangeable players and a Lipschitz-continuous potential function, Shah and Shin (2010) shows a convergence time exponential in $A$ and $1/\epsilon$ and linear in $N$, where $A$ is the number of actions per player. The latter result was extended in (Borowski and Marden, 2015) to semi-anonymous potential games, which are potential games with groups of interchangeable players.
Although finite-time convergence of log-linear learning to an $\epsilon$-efficient NE was established for different sub-classes of potential games, such guarantees do not exist for general potential games. In this paper, we focus on deriving the first finite-time convergence guarantees of log-linear learning that hold for general potential games; are polynomial in $1/\epsilon$; and remain valid under relaxed assumptions such as limited access to feedback and noisy utilities. [^2] For general potential games, an exponential dependence on $N$ in the convergence time is unavoidable since finding an $\epsilon$-efficient NE is equivalent to computing an $\epsilon$-optimizer of the potential function, which is NP-complete (Burer and Letchford, 2012).
@@ 1.2 Contributions
We study the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium in general potential games. Our contributions are summarized as follows:
  • We prove a convergence time of $\tilde{\mathcal{O}}((A^{N}/\epsilon)^{\frac{1}{\max\{\epsilon,\Delta\}}})$ to an $\epsilon$-efficient NE (Theorem 3.1), where $\Delta$ is a problem-dependent constant. If in addition, the players are interchangeable, then an $\epsilon$-NE is reached in $\tilde{\mathcal{O}}((\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$ which in contrast to general potential games is polynomial in $N$ as well (Corollary B.7).
  • We consider two variants, binary log-linear learning (Theorem 4.1) and perturbed log-linear learning (Theorem 4.3) motivated by limited feedback and noise corrupted utilities, respectively. For these variants, we show convergence guarantees that are polynomial in $1/\epsilon$.
On the technical side, past works (Blume, 1993; Young, 1993; Marden and Shamma, 2012; Shah and Shin, 2010) established asymptotic convergence guarantees by leveraging that log-linear learning induces a Markov chain. To obtain our novel finite-time results we build on this connection and develop new Markov chain results that can be summarized as follows:
  • We derive a mixing time bound for a class of Markov chains which includes those induced by log-linear learning and binary-log linear learning. It follows from an adaptation of arguments in (Diaconis and Saloff-Coste, 1996) (Lemma 2.2) and a novel bound on the log-Sobolev constant (Lemma 3.4) that we derive for this class of Markov chains.
  • We derive a tight Lipschitz constant for the known result regarding the Lipschitz-continuity of stationary distributions of Markov chains as a function of their transition matrix (Lemma 4.2). We leverage this result to study the convergence of learning rules such as perturbed log-linear learning for which the explicit stationary distributions are unknown (Theorem 4.3).
Notations
We denote by $[N]$ the set $\{1,\ldots,N\}$. For a finite set $\mathcal{X}$, we denote by $\Delta(\mathcal{X})$ the probability simplex over $\mathcal{X}$, and by $\mathds{1}_{a\in\mathcal{X}}$ the indicator function of $\mathcal{X}$. Finally, we use the big-$\mathcal{O}$ notations $\tilde{\mathcal{O}}$ and $\tilde{\Omega}$ to hide logarithmic terms.
2 Preliminaries
2.1 Problem setup
We consider a repeated potential game with $N$ players. Every player has an action set $\mathcal{A}$ of cardinality $A<\infty$, which for simplicity we assume to be the same for all players. The utility of player $i$ is a mapping $U_{i}:\mathcal{A}^{N}\rightarrow[0,1]$ from the joint action space $\mathcal{A}^{N}$ to $[0,1]$. In a potential game, the utility functions are characterized by a potential function $\Phi:\mathcal{A}^{N}\rightarrow\mathbb{R}$ such that the following holds:
@@ 2.1 Problem setup
In this paper, we assume that an initial action profile is drawn from a distribution $\mu^{0}$. The potential game is then repeated over multiple rounds. We assume that at each round $t$, a player $i\in[N]$ is uniformly chosen and selects her action based on some learning rule while the other players remain with their previous action, i.e., $a_{-i}^{t}=a_{-i}^{t-1}$.
A common assumption in game theory is that the players are rational, i.e., that they seek to maximize their utility. Under this assumption, a natural solution concept is the pure Nash equilibrium (Nash Jr, 1950).
Definition 2.1 (Nash equilibrium).
A pure Nash equilibrium is an action profile $(a_{i}^{NE})_{i\in[N]}\in\mathcal{A}^{N}$ such that every player is playing a best response to the other players’ actions $a_{-i}^{NE}$, i.e.,
$U_{i}(a_{i},a_{-i}^{NE})\leq U_{i}(a_{i}^{NE},a_{-i}^{NE}),\quad\forall i\in[N],\forall a_{i}\in\mathcal{A},$
where we define $a_{-i}:=(a_{j})_{j\in[N]\setminus\{i\}}$ as the joint action of all players except $i$.
In a pure NE, no player can improve her utility by unilaterally changing her action. Hereafter, we simply refer to a pure Nash equilibrium as a Nash equilibrium.
In potential games, several Nash equilibria may exist as shown in the two-player potential game example below. Here, action profiles $(A,A)$ and $(B,B)$ are both Nash equilibria. However, the value of the potential function may differ for different Nash equilibria, as is the case in the example.
$\bordermatrix{&A&B&\cr A&(5,2)&(-1,-2)\!\!\!\!\!\cr B&(-5,-4)&(1,4)\!\!\!\!\!\!},\quad\text{with the potential:}\quad\bordermatrix{&A&B&\cr A&4&0\!\!\!\!\!\!\!\!\cr B&-6&2\!\!\!\!\!\!\!\!}.$
@@
In this work, we are interested in learning rules that converge, in expectation, to an $\epsilon$-efficient NE, i.e., an action profile maximizing the potential up to an additive constant:
$\mathbb{E}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon,$
where the randomness stems from the distribution of action profiles at round $t$. The number of rounds needed to find an $\epsilon$-efficient NE denotes the convergence time.
Connection to Markov chains:
If all players apply a learning rule that relies exclusively on the utility of the current action profile of the other players then the learning dynamics induce a Markov chain over the state space $\mathcal{A}^{N}$. In particular, the state of the Markov chain at time $t$ corresponds to the action profile at round $t$ and the learning rule of each player specifies the transition dynamics. It follows that we can analyze the convergence time of log-linear learning by studying the associated Markov chain. We, therefore, provide relevant concepts for Markov chains in the next section.
2.2 Background on Markov chains
We briefly review concepts and properties of Markov chains used throughout this paper. Consider a time-homogeneous Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ over the state space $\mathcal{A}^{N}$ with a transition matrix $P\in\mathbb{R}^{{A^{N}}\times A^{N}}$. The ergodic theorem (Levin and Peres, 2017) states that if a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ is irreducible and aperiodic, then it has a unique stationary distribution $\mu$, and from any initial distribution $\mu^{0}$ the distribution $\mu^{t}=\mu^{0}P^{t}$ converges to $\mu$. The convergence time to the stationary distribution is quantified by the mixing time:
$t_{\operatorname{mix}}^{P}(\epsilon):=\min\{t\in\mathbb{N}\>|\>\|\mu^{t}-\mu\|_{TV}\leq\epsilon\},$ (1)
where the total variation distance is defined as $\|\mu^{t}-\mu\|_{TV}:=\frac{1}{2}\sum_{a\in\mathcal{A}^{N}}|\mu^{t}(a)-\mu(a)|$. Based on a remark in (Diaconis and Saloff-Coste, 1996, Section 3) we now derive a bound on the mixing time of Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$.
Lemma 2.2.
If $P$ is irreducible and aperiodic, then the mixing time has the following upper bound:
[…]
We briefly discuss the mixing time bound above and provide a proof and a thorough discussion in Appendix A. While classical approaches commonly bound the mixing time by the spectral gap defined in Equation (11) in Appendix A, bounds using log-Sobolev constants are often significantly tighter. Indeed, the mixing time upper bound using the log-Sobolev constant grows as $\mathcal{O}(\log\log(1/\mu_{\min}))$ whereas bounds using the spectral gap grow as $\mathcal{O}(\log(1/\sqrt{\mu_{\min}}))$. However, unlike the spectral gap deriving log-Sobolev constants can be extremely difficult. So constants have not been well-explored. In the next section in Lemma 3.4, we derive a novel bound on the log-Sobolev constant of a class of Markov chains.
@@ 3 Convergence of log-linear learning
In this section, we review the well-established log-linear learning rule (Blume, 1993) and state our main theoretical result on the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium. Our convergence analysis relies on the mixing time bound from Lemma 2.2 of the previous section.
@@ 3.1 Algorithm and background
We assume that all players follow the log-linear learning rule which is repeated over several rounds. At round $t$ a player denoted by $i$ is randomly chosen among all players and allowed to alter her action while the other players repeat their current action, i.e., $a_{-i}^{t}=a_{-i}^{t-1}$. Given full-information feedback, player $i$ observes her utility for all actions $a_{i}\in\mathcal{A}$ given the other players’ actions $a_{-i}^{t-1}$. Then, player $i$ samples an action from her strategy $p_{i}^{t}\in\Delta(\mathcal{A})$ such that:
@@
[…]
Log-linear learning induces an irreducible and aperiodic Markov chain $\{X_{t}\}_{t\in\mathbb{Z}_{+}}$ with a time-reversible transition matrix $P\in\mathbb{R}^{\mathcal{A}\times\mathcal{A}}$ (Marden and Shamma, 2012) given by:
$P_{a,\tilde{a}}=\frac{1}{N}\frac{e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}_{i}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)},$ (4)
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ of log-linear learning is given by (Blume, 1993):
The above can be verified by checking the detailed balance equations given by $\mu(a)P_{a,\tilde{a}}=\mu(\tilde{a})P_{\tilde{a},a}$ for all $a,\tilde{a}\in A^{N}$. For $\beta\rightarrow\infty$, sampling an action profile $a\in\mathcal{A}^{N}$ from the stationary distribution $\mu$ returns a maximizer of the potential $\Phi(\cdot)$ with arbitrarily high probability. Thus, when all players adhere to log-linear learning with sufficiently large $\beta$, the global outcome, in the long run, will correspond to a potential maximizer, i.e., an efficient Nash equilibrium.
For sufficiently large $\beta$, it was shown that log-linear learning converges asymptotically to a potential maximizer and thus to an efficient Nash equilibrium (Blume, 1993; Young, 1993; Marden and Shamma, 2012). Except for a few works (Montanari and Saberi, 2010; Asadpour and Saberi, 2009; Shah and Shin, 2010), none of the previous works, however, have finite-time convergence guarantees to such an efficient Nash equilibrium, and (Montanari and Saberi, 2010; Asadpour and Saberi, 2009; Shah and Shin, 2010) have additional assumptions on the potential game. Thus, in the following section, we establish a bound on the convergence time of log-linear learning in general potential games.
3.2 Convergence time
We now state our main result on the convergence time of log-linear learning to an $\epsilon$-efficient NE. Before stating the result we briefly introduce some notation. Denote by $a^{*}$ a potential maximizer, i.e., $a^{*}\in\arg\max_{a\in\mathcal{A}^{N}}\Phi(a)$. The set of $\epsilon$-optimal action profiles is defined as $\mathcal{A}^{N}(\epsilon):=\{a\in\mathcal{A}^{N}|\Phi(a)\geq\Phi(a^{*})-\epsilon\}$ with cardinality $A^{N}(\epsilon)=|\mathcal{A}^{N}(\epsilon)|$. Furthermore, the suboptimality gap is defined as $\Delta:=\min_{a\in\mathcal{A}^{N}:\Phi(a)<\Phi(a^{*})}\left(\Phi(a^{*})-\Phi(a)\right)$ and is non-negative.

### v4

@@ Abstract
This paper investigates the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium in potential games, where an efficient Nash equilibrium is defined as the maximizer of the potential function. Previous literature provides asymptotic convergence rates to efficient Nash equilibria, and existing finite-time rates are limited to potential games with further assumptions such as the interchangeability of players. We prove the first finite-time convergence to an $\epsilon$-efficient Nash equilibrium in general potential games. Our bounds depend polynomially on $1/\epsilon$, an improvement over previous bounds for subclasses of potential games that are exponential in $1/\epsilon$. We then strengthen our convergence result in two directions: first, we show that a variant of log-linear learning requiring a constant factor less feedback on the utility per round enjoys a similar convergence time; second, we demonstrate the robustness of our convergence guarantee if log-linear learning is subject to small perturbations such as alterations in the learning rule or noise-corrupted utilities.
Keywords: Efficient Nash equilibrium, game theory, log-linear learning, potential games.
@@ 1 Introduction
Interactions of multiple agents are at the heart of many applications, including transportation networks, auctions, telecommunication networks, and multi-robot systems. A common solution concept to describe the outcomes of multi-agent systems is the Nash equilibrium [33].
Thus, a natural question is whether strategic players can learn a Nash equilibrium and, if so, at what speed they can learn it. As games can have multiple Nash equilibria of different quality, in terms of social welfare, for example, it is important to understand which Nash equilibrium is learned.
A class of games that are suitable for learning are potential games [29], where joint actions that maximize the potential function correspond to Nash equilibria. If social welfare and the potential function are aligned in the sense that an increase in social welfare is associated with an increase in potential [35], then a Nash equilibrium that maximizes the potential also maximizes the social welfare. This is the case for identical interest games and distributed welfare games [27], where social welfare is given by the aggregated players’ utilities.
In this paper, motivated by the connection between the potential function and social welfare, we study the speed of convergence of a decentralized learning algorithm to an approximately efficient Nash equilibrium. This question is important, as it determines how quickly desirable outcomes can be achieved through decentralized learning.
@@ 1.1 Related work
In potential games, many learning rules were shown to converge to an arbitrary Nash equilibrium, such as iterative best-response dynamics [37, 8, 15], no-regret algorithms [21, 36, 18], and fictitious play [28, 29]. On the other hand, log-linear learning [10, 40] is the only known algorithm to converge to a specific Nash equilibrium, namely, the potential maximizer. Past works provide asymptotic convergence guarantees [10, 40] as well as an asymptotic rate of convergence to a potential function maximizer [39]. Such asymptotic convergence guarantees were also shown for slight variations of log-linear learning which include synchronous updates [26], utilities that are corrupted by noise [22], or payoff-based or two points of feedback per round [26, 1] Another line of research studied the mixing time of the Markov chain induced by log-linear learning [7, 5] and its transient behavior prior to reaching the stationary distribution [6], but did not explicitly relate the stationary distribution to the set of efficient Nash equilibria.
While asymptotic convergence and convergence rates characterize the long-run behavior and the asymptotic speed, respectively, at which log-linear learning approaches an efficient equilibrium, finite-time convergence provides explicit bounds on the number of steps required to reach an $\epsilon$-efficient Nash equilibrium, making it particularly desirable in practice.
Few past works provide finite-time guarantees for log-linear learning to an $\epsilon$-efficient Nash equilibrium, an action profile whose potential is $\epsilon$-close to the maximum value. In specific potential games, namely, in atomic routing games with polynomial costs, [2] derives a convergence time that is exponential in $1/\epsilon$ and polynomial in $N$, the number of players. Moreover, in games with a graph structure between players, [30, 31] prove a convergence time, which is exponential in $1/\epsilon$ and $N$ in the worst case. Finally, in potential games with interchangeable players and a Lipschitz-continuous potential function, [38] shows a convergence time exponential in $A$ and $1/\epsilon$ and linear in $N$, where $A$ is the number of actions per player. The latter result was extended to semi-anonymous potential games [11], which consist of groups of interchangeable players.
While finite-time convergence of log-linear learning has been proven for subclasses of potential games, to our knowledge, it is not established for general potential games.
@@ 1.2 Contributions
In this paper, we derive the first finite-time convergence guarantees of log-linear learning to an $\epsilon$-efficient Nash equilibrium in general potential games. Our contributions are:
  • We prove a convergence time of $\tilde{\mathcal{O}}((A^{N}/\epsilon)^{\frac{1}{\Delta}})$ to an $\epsilon$-efficient Nash equilibrium (Theorem 3.1), where the suboptimality gap $\Delta$ is a problem-dependent constant.
  • If in addition, the players are interchangeable, then an $\epsilon$-Nash equilibrium is reached in $\tilde{\mathcal{O}}((\frac{N^{A}}{\epsilon})^{\frac{1}{\Delta}})$ which in contrast to general potential games is polynomial in $N$ (Corollary 3.3).
  • We consider two variants of log-linear learning: binary log-linear learning and perturbed log-linear learning motivated by limited feedback and noise corrupted utilities, respectively. We prove a convergence time of $\tilde{\mathcal{O}}((A^{N}/\epsilon)^{\frac{1}{\Delta}})$ (Theorem 5.1) and $\tilde{\mathcal{O}}((A^{N}/\epsilon)^{\frac{1}{\Delta}N(1+\xi)})$ (Theorem 5.2) , respectively.
On the technical side, past works [10, 40] established that log-linear learning induces a Markov chain. To obtain our novel finite-time results, we build on this connection and develop new results about Markov chain mixing times and stationary distributions as follows:
  • We use mixing-time bounds based on the so-called log-Sobolev constant of the Markov chain [16] to establish finite-time convergence guarantees for log-linear learning and its variants. To this end, we derive a novel bound on the log-Sobolev constant of a class of Markov chains which includes those induced by log-linear- and binary log-linear learning (Lemma 4.2).
  • We derive a tight Lipschitz constant of stationary distributions of Markov chains as a function of their transition matrix (Lemma 5.3). We leverage this result to study the convergence of perturbed log-linear learning for which the stationary distribution is unknown (Theorem 5.2).
Notations: We denote by $[N]$ the set $\{1,\ldots,N\}$. For a finite set $\mathcal{X}$, we denote by $\Delta(\mathcal{X})$ the probability simplex over $\mathcal{X}$, and by $\mathds{1}_{a\in\mathcal{X}}$ the indicator function of $\mathcal{X}$. Finally, we use the big-$\mathcal{O}$ notations $\tilde{\mathcal{O}}$ and $\tilde{\Omega}$ to hide logarithmic terms.
2 Problem setup
We consider a potential game with $N$ players. Every player has an action set $\mathcal{A}$ of cardinality $A<\infty$, which for simplicity we assume to be the same for all players. The utility of player $i$ is a mapping $U_{i}:\mathcal{A}^{N}\rightarrow[0,1]$, where $\mathcal{A}^{N}$ is the joint action space. In a potential game, the utility functions are characterized by a potential function $\Phi:\mathcal{A}^{N}\rightarrow\mathbb{R}$ such that:
@@ 2 Problem setup
A common solution concept is the Nash equilibrium [34], at which no player can improve her utility by unilaterally changing her action.
Definition 2.1.
A Nash equilibrium is an action profile $(a_{i})_{i\in[N]}\in\mathcal{A}^{N}$ that satisfies:
$U_{i}(\bar{a}_{i},a_{-i})\leq U_{i}(a_{i},a_{-i}),\quad\forall i\in[N],\forall\bar{a}_{i}\in\mathcal{A},$
where $a_{-i}:=(a_{j})_{j\in[N]\setminus\{i\}}$ is the action of all players but $i$.
Generally, a game may have several Nash equilibria, as shown in the two-player potential game below. Here, action profiles $(A,A)$ and $(B,B)$ are both Nash equilibria. However, the value of the potential function may differ for different Nash equilibria, as is the case in the example.
@@ A Nash equilibrium is an action profile $(a_{i})_{i\in[N]}\in\mathcal{A}^{N}$ that satisfies:
This example motivates the distinction between a Nash equilibrium and an efficient Nash equilibrium, defined as:
$a^{*}\in\operatorname*{arg\,max}_{a\in\mathcal{A}^{N}}\Phi(a).$ (1)
Note that such $a^{*}$ exists and is a Nash equilibrium [29].
In a game setting, players act independently and do not have knowledge of the other players’ utilities, nor do they share their utility with a central authority. As a result, it is impossible to enumerate over the set of joint actions to identify the potential function and thus find its maximizer.
Instead, in this work, we consider a repeated game setting, where the game unfolds over multiple rounds. Our focus is on learning rules that converge to an $\epsilon$-efficient Nash equilibrium.
Definition 2.2.
An action profile $a^{*}\in\mathcal{A}^{N}$ is an $\epsilon$-efficient Nash equilibrium if it satisfies $\Phi(a^{*})\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon$ for $\epsilon\in(0,1)$.
@@ 3 Convergence of log-linear learning
In this section, we introduce the well-established log-linear learning rule [10] and state our main result on the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium.
@@ 3.1 Algorithm and background
We consider a repeated game setting in which all players follow log-linear learning. In the initial round, players initialize their action randomly according to some distribution $\mu^{0}$. Thereafter, at round $t$, a player denoted by $i$ is randomly chosen among all players and allowed to alter her action while the other players repeat their current action, i.e., $a_{-i}^{t}=a_{-i}^{t-1}$. Player $i$ observes her utility for all actions $a_{i}\in\mathcal{A}$ given the other players’ actions $a_{-i}^{t-1}$. Then, player $i$ samples an action from her strategy $p_{i}^{t}\in\Delta(\mathcal{A})$ such that:
@@
where parameter $\beta$ measures rationality: for large $\beta$ player $i$ likely selects a best response $a_{i}^{t}\in\operatorname*{arg\,max}_{a_{i}\in\mathcal{A}}U_{i}(a_{i},a_{-i}^{t-1})$; and for $\beta=0$ player $i$ samples $a_{i}^{t}$ uniformly.
Log-linear learning induces an irreducible and aperiodic Markov chain $\{X_{t}\}_{t\in\mathbb{Z}_{+}}$ over state space $\mathcal{A}^{N}$ with a time-reversible transition matrix $P\in\mathbb{R}^{A^{N}\times A^{N}}$ [26] given by:
$P_{a,\tilde{a}}=\frac{1}{N}\frac{e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}.$ (3)
Here $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$ is the set of action profiles $\tilde{a}\in\mathcal{A}^{N}$ that differ from action profile $a\in\mathcal{A}^{N}$ in at most one player’s action. The stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ of log-linear learning is given by [10]:
$\mu(a)=\frac{e^{\beta\Phi(a)}}{\sum_{\tilde{a}\in\mathcal{A}^{N}}e^{\beta\Phi(\tilde{a})}},\quad\forall a\in\mathcal{A}^{N}.$ (4)
The above can be verified by checking the detailed balance equations corresponding to $\mu$ [32]. [^1] Detailed balance holds if $\mu(a)P_{a,\tilde{a}}=\mu(\tilde{a})P_{\tilde{a},a}$ for all $a,\tilde{a}\in A^{N}$. It follows that we can analyze the convergence time of log-linear learning by studying the associated Markov chain.
Previous works [10, 40, 26] show that for sufficiently large $\beta$ log-linear learning converges asymptotically to a potential function maximizer and thus to an efficient Nash equilibrium. With the exception of a few works [31, 2, 38] which make additional assumptions on the potential game, none of the previous works, however, provide finite-time convergence guarantees. Thus, in the following section, we establish our main result on the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium in general potential games.
3.2 Convergence time in general potential games
We first introduce some notation needed to state our main result. Denote by $a^{*}$ a potential maximizer, namely $a^{*}\in\arg\max_{a\in\mathcal{A}^{N}}\Phi(a)$ and by $\mathcal{A}_{*}^{N}:=\{a^{*}\in\mathcal{A}^{N}|a^{*}\in\arg\max_{a\in\mathcal{A}^{N}}\Phi(a)\}$ the set of optimal action profiles with cardinality $A_{*}^{N}=|\mathcal{A}_{*}^{N}|$. Moreover, we define the suboptimality gap as
$\Delta:=\min_{a\in\mathcal{A}_{P}:\Phi(a)<\Phi(a^{*})}\bigl(\Phi(a^{*})-\Phi(a)\bigr).$
By construction, $\Delta\geq 0$. The degenerate case $\Delta=0$ is trivial, since it implies that all action profiles achieve the optimal potential value and hence are efficient Nash equilibria. Consequently, throughout the remainder of this paper, we assume $\Delta>0$.